OpenStudy (boxman61):

Answer Check- A drug is dissolved in a solution at a concentration of 350 mg drug/10. cm3 solution. The drug dosage is 15 mg per kilogram of body weight. Show on one line how many mL of the drug solution is required for a 180 lb person.

OpenStudy (boxman61):

1lb=454g 180*454= 81,720 grams or 81.72Kg 81.72*15= 1225.8 mg drug dosage required 1225.8/350=3.5 (drug/solution required) 3.5*10.cm^3 = 35cm^3 of solution or 35mL of solution

OpenStudy (welshfella):

yea That is correct

OpenStudy (boxman61):

thank you!!

OpenStudy (welshfella):

yw